OpenStudy (anonymous):

Solve the system by substitution. x - 3y = 42 and x - 6y = 8

OpenStudy (jack1):

ENQ 1 x - 3y = 42 ENQ 2 x - 6y = 8 what does x equal if u rearrange only eqn 2?

OpenStudy (anonymous):

do not just give the answer...jack is trying to explain it

OpenStudy (anonymous):

im sorry i did it wrong it x - 3y= 4 2x - 6y = 8

OpenStudy (anonymous):

subtract those two equation and you will get the value of y then put value of y in any equation to get value of x...

OpenStudy (anonymous):

so much for explaining it

OpenStudy (jack1):

^^^ or that ;D

OpenStudy (anonymous):

those two lines are coincident so there will be infinite solutions

OpenStudy (anonymous):

thanks ^

OpenStudy (anonymous):

no prob.. (:

OpenStudy (anonymous):

6x + 3y = -6 2x + y = -2

OpenStudy (anonymous):

what about that equation?

OpenStudy (anonymous):

still equations are coincident

OpenStudy (jack1):

EQN 1 6x + 3y = -6 EQN 2 2x + y = -2 put equation 2 in terms of y = ________ ? (ie rearrange the equation until you get y = something x) then substitute that y value into the other equation

OpenStudy (anonymous):

thanks both of yall